Three students were given a differential equation y''+π ^2y=0 to solve. 


1
Expert's answer
2021-05-26T03:16:35-0400

"y''+\pi^2y=0"

"\dfrac{d^2y}{dx^2}=y''=D^2y"

"D^2y+\pi^2y=0"

"(D^2+\pi^2)y=0"


Auxiliary equation,

"D^2+\pi^2=0"

"D^2=-\pi^2"

"D=\pm i\pi"


Complimentary function is given by,

"y=C_1e^{i\pi}+C_2e^{-i\pi}"